D Question 10 2.6 pts Suppose the demand and supply curves for a particular product are given below. D P160- 2Q s: P-40+ q What is the equilibrium price, P, and equilibrium quantity, Q? None of the other options. P = 47.66; Q-5667. P=90; Q = 50. P 60; Q 50. P80:Q-40please answer
Answer #1

At equilibrium demand is equal to supply

160-2Q = 40+Q

120= 3Q

Q= 40

P= 40+40= 80

P= 80; Q= 40